# signature_mor.bib

### v1

@@ 3.1 Identifying less relevant signature information
We introduce the fundamental solution to the signature equation (3.1) as the process $\Phi(t,t_{0})$ satisfying
(3.2) $\mathrm{d}\Phi(t,t_{0})=A\Phi(t,t_{0})\mathrm{d}t+\sum_{i=2}^{d}N_{i}\Phi(t,t_{0})\mathrm{d}B_{t}^{i},\quad\Phi(t_{0},t_{0})=I\in\mathbb{R}^{n\times n},$
with $t_{0}\in[0,T)$. Therefore, we can write $\mathcal{X}_{t}=\Phi(t,t_{0})\mathcal{X}_{t_{0}}$ for all $t\in[t_{0},T]$. Based on that, we set
@@
We can choose an orthonormal basis (ONB) consisting of eigenvectors of these time-averaged covariances to represent the truncated signature. Let $(p_{k})_{k=1,\dots,n}$ be an ONB for $\mathbb{R}^{n}$ of eigenvectors of $P$ and $(q_{k})_{k=1,\dots,n}$ the one associated to $Q$. This yields the following
(3.5) $\mathcal{X}_{t}$ $=\sum_{k=1}^{n}\langle\mathcal{X}_{t},p_{k}\rangle_{\mathbb{R}^{n}}\,p_{k},\quad t\in[0,T],$
$\mathcal{Y}_{t}$ $=L\Phi(t,t_{0})\mathcal{X}_{t_{0}}=\sum_{k=1}^{n}\langle\mathcal{X}_{t_{0}},q_{k}\rangle_{\mathbb{R}^{n}}\,L\Phi(t,t_{0})q_{k},\quad t\in[t_{0},T].$
The next proposition tells us which directions $p_{k}$ and $q_{k}$ are of minor relevance.
@@
Let $(\lambda_{k})_{k=1,\dots,n}$ and $(\mu_{k})_{k=1,\dots,n}$ be the eigenvalues of the ONB $(p_{k})_{k=1,\dots,n}$ and $(q_{k})_{k=1,\dots,n}$, respectively. Then, we obtain for the $k$-th summands of (3.5) that
$\mathbb{E}\int_{0}^{T}\langle\mathcal{X}_{u},p_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathrm{d}u$ $=\lambda_{k},$
$\mathbb{E}\int_{t_{0}}^{T}\langle\mathcal{X}_{t_{0}},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\left\|L\Phi(u,t_{0})q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$ $\leq\mu_{k}\mathbb{E}\langle\mathcal{X}_{t_{0}},q_{k}\rangle_{\mathbb{R}^{n}}^{2}.$
@@
The first identity is trivial as $\mathbb{E}\int_{0}^{T}\langle\mathcal{X}_{u},p_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathrm{d}u=p_{k}^{\top}\int_{0}^{T}\mathbb{E}[\mathcal{X}_{u}\mathcal{X}_{u}^{\top}]\mathrm{d}u\,p_{k}=p_{k}^{\top}Pp_{k}=\lambda_{k}$ exploiting that $p_{k}$ has norm $1$. We further note that $\mathcal{X}_{t_{0}}$ is $\mathcal{F}_{t_{0}}$-measurable and that $u\mapsto\Phi(u,t_{0})$ only depends on increments of $B$ after $t_{0}$. Therefore, they are independent and hence $\mathbb{E}[\langle\mathcal{X}_{t_{0}},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\left\|L\Phi(u,t_{0})q_{k}\right\|_{\mathbb{R}^{p}}^{2}]=\mathbb{E}\langle\mathcal{X}_{t_{0}},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathbb{E}\left\|L\Phi(u,t_{0})q_{k}\right\|_{\mathbb{R}^{p}}^{2}$. Now, we obtain by substitution that
$\mathbb{E}\int_{t_{0}}^{T}\left\|L\Phi(u,t_{0})q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$ $=\mathbb{E}\int_{t_{0}}^{T}\left\|L\Phi(u-t_{0},0)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$
$=\int_{0}^{T-t_{0}}\mathbb{E}\left\|L\Phi(u,0)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u\leq q_{k}Qq_{k}=\mu_{k}.$
Above, we exploited that $\mathbb{E}\big{[}\Phi(u,t_{0})q_{k}q_{k}^{\top}\Phi(u,t_{0})^{\top}\big{]}=\mathbb{E}\big{[}\Phi(u-t_{0},0)q_{k}q_{k}^{\top}\Phi(u-t_{0},0)^{\top}\big{]}$ as both expressions satisfy (3.6) with $M=q_{k}q_{k}^{\top}$. This concludes the proof. ∎
Proposition 3.3 tell us that $p_{k}$ has a low contribution to $\mathcal{X}_{t}$ if the associated eigenvalue $\lambda_{k}$ is small. Moreover, neglecting $q_{k}$ in $\mathcal{X}_{t_{0}}$ has a minor impact on the quantity of interest after $t_{0}$ if $\mu_{k}$ is small unless $\mathcal{X}_{t_{0}}$ is very large in the direction $q_{k}$. Consequently, we know that we can remove directions $p_{k}$ and $q_{k}$ that correspond to small eigenvalues. In order to be able to do this simultaneously, we construct a coordinate transform below that ensures that $p_{k}=q_{k}=\operatorname{e}_{k}$. Here, $\operatorname{e}_{k}$ is $k$-th canonical basis vector of $\mathbb{R}^{n}$. Therefore, unimportant direction can be identified with components of a transformed signature. Before stressing this aspect further, a strategy for the computation of $P$ and $Q$ is provided in the following section.
@@ 3.2 Computation of $P$ and $Q$
The later dimension reduction procedure relies on having $P$ and $Q$ available. For that reason, let us briefly discuss how these matrices can be computed in practice.
Let us introduce the Lyapunov operator $\mathcal{L}(Z)=AZ+ZA^{\top}+\sum_{i,j=2}^{d}N_{i}ZN_{j}^{\top}k_{ij}$, where $Z$ is a matrix of suitable dimension. This notion is used to recall the following well-known result.
@@
Given $M\in\mathbb{R}^{n\times n}$, the function $t\mapsto\mathbb{E}\big{[}\Phi(t,t_{0})M\Phi(t,t_{0})^{\top}\big{]}$ solves
(3.6) $\frac{\mathrm{d}}{\mathrm{d}t}Z_{t}=\mathcal{L}\left(Z_{t}\right),\quad Z_{t_{0}}=M,\quad t\geq t_{0}.$
@@
We define $\Phi_{t}:=\Phi(t,t_{0})$ and make use of Itô’s product rule yielding
@@
Taking the expectation, we obtain $\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}\big{[}\Phi_{t}M\Phi_{t}^{\top}\big{]}=\mathcal{L}\left(\mathbb{E}\big{[}\Phi_{t}M\Phi_{t}^{\top}\big{]}\right).$
Let $\operatorname{vec}(\cdot)$ be the vectorization of a matrix. Using its relation to the Kronecker product $\tilde{\otimes}$ between two matrices, (3.6) becomes
(3.7) $\frac{\mathrm{d}}{\mathrm{d}t}\operatorname{vec}(Z_{t})=\mathcal{K}\operatorname{vec}(Z_{t}),\quad\operatorname{vec}(Z_{t_{0}})=\operatorname{vec}(M),\quad t\geq t_{0},$
where the matrix representation of the Lyapunov operator is
@@
As $P$ defined in (3.3) is the integral of the solution of (3.6) with $t_{0}=0$ and $M=zz^{\top}$, we therefore have that
@@
inserting the solution representation for (3.7). According to Remark 2.14, the matrices $N_{1},\dots,N_{d}$ are nilpotent with order $m+1$ leading to $\mathcal{K}^{j}=0$ for $j\geq 2m+1$ and hence $\operatorname{e}^{\mathcal{K}t}=\sum_{j=0}^{2m}\frac{t^{j}}{j!}\mathcal{K}^{j}$. Exploiting this fact yields
@@
$\operatorname{e}^{\mathcal{K}(t-t_{0})}\operatorname{vec}(M)=\operatorname{vec}\Big{(}\mathbb{E}\Big{[}\Phi(t,t_{0})M\Phi(t,t_{0})^{\top}\Big{]}\Big{)}=\mathbb{E}\Big{[}\Phi(t,t_{0})\,\tilde{\otimes}\,\Phi(t,t_{0})\Big{]}\operatorname{vec}(M).$
Since this identity is true for all matrices $M$, we find that $\mathbb{E}\Big{[}\Phi(t,t_{0})\,\tilde{\otimes}\,\Phi(t,t_{0})\Big{]}=\operatorname{e}^{\mathcal{K}(t-t_{0})}$ and consequently
@@ 3.3 Reduced order signature approximation
We conduct a state space transformation via a nonsingular matrix $\mathcal{T}\in\mathbb{R}^{n\times n}$. To do so, we define $\hat{\mathcal{X}}_{t}=\mathcal{T}\mathcal{X}_{t}$ and and associated coefficients $(\hat{A},\hat{N}_{i},\hat{L})=(\mathcal{T}A\mathcal{T}^{-1},\mathcal{T}N_{i}\mathcal{T}^{-1},L\mathcal{T}^{-1})$. The purpose of this transformation is that an equivalent system (same quantity of interest) is supposed to be obtained, in which the redundant information can be removed easily by truncation of state components. In particular, $\mathcal{T}$ is chosen in a way that it simultaneously diagonalizes $P$ and $Q$. Hence, the corresponding eigenvectors $(p_{k})$ and $(q_{k})$ are the canonical basis $(\operatorname{e}_{k})$ of $\mathbb{R}^{n}$ at the same time. Exploiting Proposition 3.3, we can then identify unimportant components $\langle\hat{\mathcal{X}}_{\cdot},\operatorname{e}_{k}\rangle_{\mathbb{R}^{n}}$ of $\hat{\mathcal{X}}$ and truncate those. Let us refer to the discussion below this proposition once more and provide further details below.
@@
meaning that $\mathcal{Y}$ is invariant under this transformation. Now, we see directly from (3.2) that the fundamental solution of (3.12) is $\hat{\Phi}(t,t_{0})=\mathcal{T}\Phi(t,t_{0})\mathcal{T}^{-1}$. Consequently, the time-averaged covariances of (3.12) are
(3.13) $\hat{P}$ $:=\int_{0}^{T}\mathbb{E}\left[\hat{\Phi}(u,0)\mathcal{T}z(\mathcal{T}z)^{\top}\hat{\Phi}(u,0)^{\top}\right]\mathrm{d}u=\mathcal{T}P\mathcal{T}^{\top},$
$\hat{Q}$ $:=\int_{0}^{T}\mathbb{E}\left[\hat{\Phi}(u,0)^{\top}\hat{L}^{\top}\hat{L}\hat{\Phi}(u,0)\right]\mathrm{d}u=\mathcal{T}^{-\top}Q\mathcal{T}^{-1}.$
The following proposition states the particular transformation required for a simultaneous diagonalization of the time-averaged covariances.
@@
Given that we have used the balancing transformation (3.14) in (3.12), we can identify less relevant directions with components of $\hat{\mathcal{X}}$ that are associated to small eigenvalues $\sigma_{k}^{2}$ of $PQ$. This is a consequence of Proposition 3.3, where the transformation ensures that $p_{k}=q_{k}$ is the $k$-th column of the $n\times n$ identity matrix and $\lambda_{k}=\mu_{k}=\sigma_{k}$. For that reason, the spectrum of $PQ$ delivers a good truncation criterion and therefore the intuition on how to fix the reduced dimension $\widetilde{n}\ll n$. In order to find the reduced equation, we partition the solution of (3.12) with $\mathcal{T}$ like in (3.14) as follows $\hat{\mathcal{X}}_{t}=\begin{bmatrix}\hat{\mathcal{X}}_{t}^{1}\\ \hat{\mathcal{X}}_{t}^{2}\end{bmatrix}$. $\hat{\mathcal{X}}_{t}^{1}$ taking values in $\mathbb{R}^{\widetilde{n}}$ corresponds to the large values $\sigma_{1},\ldots,\sigma_{\widetilde{n}}$ and $\hat{\mathcal{X}}_{t}^{2}$ to the small values $\sigma_{\widetilde{n}+1},\ldots,\sigma_{n}$ in the sense that $\sigma_{\widetilde{n}+1}\ll\sigma_{\widetilde{n}}$. The respective partition of the balanced coefficients

### v2

@@ 3.1 Identifying less relevant signature information
The aim of the MOR procedure proposed in this paper is to find a low dimension process $\widetilde{\mathcal{X}}$ that satisfies a linear SDE and approximates (3.1) in the sense that a linear map of $\widetilde{\mathcal{X}}$ can accurately reproduce $\mathcal{Y}$ which usually is a linear map of the signature. Therefore, we derive computable algebraic objects that identify unimportant information in the SDE solved by $\mathcal{X}$} and further detect redundancies in the linear map $L{\mathcal{X}}$ determined by the matrix $L$. We introduce the fundamental solution of the signature equation (3.1) in this context. It is the process $\Phi(t,s)$ satisfying
(3.2) $\mathrm{d}\Phi(t,s)=A\Phi(t,s)\mathrm{d}t+\sum_{i=2}^{d}N_{i}\Phi(t,s)\mathrm{d}B_{t}^{i},\quad\Phi(s,s)=I\in\mathbb{R}^{n\times n},$
with $s\in[0,T)$. Therefore, we can write $\mathcal{X}_{t}=\Phi(t,s)\mathcal{X}_{s}$ for all $t\in[s,T]$. Based on the fundamental solution, we set
@@
$P$ is defined to detect dominant subspaces in the SDE in (3.1), whereas $Q$ is needed to identify which state directions are not important in $\mathcal{Y}=L\mathcal{X}$. For that reason, $P$ involves the initial state $z$ of the SDE and $Q$ contains the matrix $L$ that represents a linear map of the signature. In this context two representations of $\mathcal{X}$ are required using different orthonormal bases (ONB) of $\mathbb{R}^{n}$. In particular, we choose ONB consisting of eigenvectors of $P$ and $Q$ to represent the truncated signature. Let $(p_{k})_{k=1,\dots,n}$ be an ONB for $\mathbb{R}^{n}$ of eigenvectors of $P$ and $(q_{k})_{k=1,\dots,n}$ the one associated with $Q$. This yields the following
(3.5) $\mathcal{X}_{t}=\sum_{k=1}^{n}\langle\mathcal{X}_{t},p_{k}\rangle_{\mathbb{R}^{n}}\,p_{k},\quad\mathcal{X}_{t}=\sum_{k=1}^{n}\langle\mathcal{X}_{t},q_{k}\rangle_{\mathbb{R}^{n}}\,q_{k},\quad t\in[0,T],$
First, we find the directions $p_{k}$, in which $\mathcal{X}$ is small, since those are not relevant. In order to do so, we look at the associated coefficient $\langle\mathcal{X},p_{k}\rangle_{\mathbb{R}^{n}}$ in the following proposition.
@@
Let $(\lambda_{k})_{k=1,\dots,n}$ be the eigenvalues corresponding to the ONB $(p_{k})_{k=1,\dots,n}$. Then, we have that
$\mathbb{E}\int_{0}^{T}\langle\mathcal{X}_{u},p_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathrm{d}u$ $=\lambda_{k}.$
@@
This identity is trivial as $\mathbb{E}\int_{0}^{T}\langle\mathcal{X}_{u},p_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathrm{d}u=p_{k}^{\top}\int_{0}^{T}\mathbb{E}[\mathcal{X}_{u}\mathcal{X}_{u}^{\top}]\mathrm{d}u\,p_{k}=p_{k}^{\top}Pp_{k}=\lambda_{k}$ exploiting that $p_{k}$ has norm $1$. ∎
Second, we use (3.5) to answer which directions $q_{k}$ do not matter for $\mathcal{Y}=L\mathcal{X}$. To be more precise, we look at $\mathcal{X}_{s}$ for each $s\in[0,T)$ and investigate the impact on $\mathcal{Y}$ on the interval $[s,T]$ when omitting a direction $q_{k}$. In fact, we study the following difference
(3.6) $\mathcal{Y}_{t}-L\Phi(t,s)\sum_{j=1\atop j\neq k}^{n}\langle\mathcal{X}_{s},q_{j}\rangle_{\mathbb{R}^{n}}\,q_{j}=\langle\mathcal{X}_{s},q_{k}\rangle_{\mathbb{R}^{n}}\,L\Phi(t,s)q_{k},\quad t\in[s,T],$
for $k\in\{1,\dots,n\}$ using the second representation in (3.5) which yields that $\mathcal{Y}_{t}=L\Phi(t,s)\mathcal{X}_{s}=\sum_{j=1}^{n}\langle\mathcal{X}_{s},q_{j}\rangle_{\mathbb{R}^{n}}\,L\Phi(t,s)q_{j}$. In the following proposition, the $L^{2}$-norm of the difference in (3.6) is investigated. If it is small, it means that $q_{k}$ plays a minor role.
Proposition 3.5.
Let $(\mu_{k})_{k=1,\dots,n}$ be the eigenvalues corresponding to the ONB $(q_{k})_{k=1,\dots,n}$. Then, we obtain that
$\mathbb{E}\int_{s}^{T}\langle\mathcal{X}_{s},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\left\|L\Phi(u,s)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$ $\leq\mu_{k}\mathbb{E}\langle\mathcal{X}_{s},q_{k}\rangle_{\mathbb{R}^{n}}^{2}$
for any $s\in[0,T)$.
Proof.
We note that $\mathcal{X}_{s}$ is $\mathcal{F}_{s}$-measurable and that $u\mapsto\Phi(u,s)$ only depends on increments of $B$ after time $s$. Therefore, they are independent and hence we obtain $\mathbb{E}[\langle\mathcal{X}_{s},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\left\|L\Phi(u,s)q_{k}\right\|_{\mathbb{R}^{p}}^{2}]=\mathbb{E}\langle\mathcal{X}_{s},q_{k}\rangle_{\mathbb{R}^{n}}^{2}\mathbb{E}\left\|L\Phi(u,s)q_{k}\right\|_{\mathbb{R}^{p}}^{2}$. Now, we obtain by substitution that
$\mathbb{E}\int_{s}^{T}\left\|L\Phi(u,s)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$ $=\mathbb{E}\int_{s}^{T}\left\|L\Phi(u-s,0)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u$
$=\int_{0}^{T-s}\mathbb{E}\left\|L\Phi(u,0)q_{k}\right\|_{\mathbb{R}^{p}}^{2}\mathrm{d}u\leq q_{k}Qq_{k}=\mu_{k}.$
Above, we exploited that $\mathbb{E}\big[\Phi(u,s)q_{k}q_{k}^{\top}\Phi(u,s)^{\top}\big]=\mathbb{E}\big[\Phi(u-s,0)q_{k}q_{k}^{\top}\Phi(u-s,0)^{\top}\big]$ as both expressions satisfy (3.7) with $M=q_{k}q_{k}^{\top}$. This concludes the proof. ∎
Proposition 3.4 tells us that $p_{k}$ has a low contribution to $\mathcal{X}_{t}$ if the associated eigenvalue $\lambda_{k}$ is small. Moreover, Proposition 3.5 indicates that neglecting $q_{k}$ in $\mathcal{X}_{s}$ has a minor impact on the quantity of interest $\mathcal{Y}$ after time $s$ if $\mu_{k}$ is small (unless $\mathcal{X}_{s}$ is very large in the direction $q_{k}$). Consequently, we know that we can remove directions $p_{k}$ and $q_{k}$ that correspond to small eigenvalues. In order to be able to do this simultaneously, we construct a coordinate transform below that ensures that $p_{k}=q_{k}=\operatorname{e}_{k}$. Here, $\operatorname{e}_{k}$ is $k$-th canonical basis vector of $\mathbb{R}^{n}$. Therefore, unimportant directions can be identified with components of a transformed signature. Before stressing this aspect further, a strategy for the computation of $P$ and $Q$ is provided in the following section.
@@ 3.2 Computation of $P$ and $Q$
Computability of the reduced system is a crucial aspect of our approach. The later dimension reduction procedure relies on having $P$ and $Q$, defined in (3.3) and (3.4), available. For that reason, let us briefly discuss how these matrices can be computed in practice. We introduce the Lyapunov operator
$\mathcal{L}(Z)=AZ+ZA^{\top}+\sum_{i,j=2}^{d}N_{i}ZN_{j}^{\top}k_{ij},$
where $Z$ is a matrix of suitable dimension. $P$ and $Q$ will be computed from equations involving this operator (or its adjoint). By definition in (3.3), $P$ is the integral of an expectation. Therefore, we recall the following well-known result that states that this expectation solves a matrix ordinary differential equation.
@@
Given $M\in\mathbb{R}^{n\times n}$, the function $t\mapsto\mathbb{E}\big[\Phi(t,s)M\Phi(t,s)^{\top}\big]$ solves
(3.7) $\frac{\mathrm{d}}{\mathrm{d}t}Z_{t}=\mathcal{L}\left(Z_{t}\right),\quad Z_{s}=M,\quad t\geq s.$
@@
We define $\Phi_{t}:=\Phi(t,s)$ and make use of Itô’s product rule yielding
@@
Taking the expectation, we obtain $\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}\big[\Phi_{t}M\Phi_{t}^{\top}\big]=\mathcal{L}\left(\mathbb{E}\big[\Phi_{t}M\Phi_{t}^{\top}\big]\right)$ using the Fubini theorem and the fact that the Itô integral has zero mean. Both properties hold automatically in the Brownian setting. However, in a general semimartingale framework, additional assumptions are required to justify their use.
Let $\operatorname{vec}(M)=\big(\begin{matrix}\mathfrak{m}_{1}^{\top}&\mathfrak{m}_{2}^{\top}&\dots&\mathfrak{m}_{n}^{\top}\end{matrix}\big)^{\top}$ be the vectorization of a matrix $M$ with $n$ columns $\mathfrak{m}_{1},\dots,\mathfrak{m}_{n}$. Using its relation to the Kronecker product $\tilde{\otimes}$ between two matrices, (3.7) becomes
(3.8) $\frac{\mathrm{d}}{\mathrm{d}t}\operatorname{vec}(Z_{t})=\mathcal{K}\operatorname{vec}(Z_{t}),\quad\operatorname{vec}(Z_{s})=\operatorname{vec}(M),\quad t\geq s,$
where the matrix representation of the Lyapunov operator $\mathcal{L}$ is
@@
As $P$ defined in (3.3) is the integral of the solution of (3.7) with $s=0$ and $M=zz^{\top}$, we therefore have that
@@
inserting the solution representation for (3.8). According to Remark 2.1, the matrices $N_{1},\dots,N_{d}$ are nilpotent with order $m+1$ leading to $\mathcal{K}^{j}=0$ for $j\geq 2m+1$ and hence $\operatorname{e}^{\mathcal{K}t}=\sum_{j=0}^{2m}\frac{t^{j}}{j!}\mathcal{K}^{j}$.
Remark 3.7.
The degree of nilpotency of $\mathcal{K}$ can be argued as follows. It is determined by $A\,\tilde{\otimes}\,I+I\,\tilde{\otimes}\,A$, since this term involves the non-nilpotent identity matrix. As $A\,\tilde{\otimes}\,I$ and $I\,\tilde{\otimes}\,A$ commute, we find that
$\big(A\,\tilde{\otimes}\,I+I\,\tilde{\otimes}\,A\big)^{n}=\sum_{k=0}^{n}\binom{n}{k}\big(A\,\tilde{\otimes}\,I\big)^{k}\big(I\,\tilde{\otimes}\,A\big)^{n-k}=\sum_{k=0}^{n}\binom{n}{k}A^{k}\,\tilde{\otimes}\,A^{n-k}.$
Using Remark 2.1, we know that $N_{1}$ and hence also $A$ is nilpotent of degree $m+1$. Therefore, given that $0\leq n\leq 2m$ there is a $k=0,1,\dots,n$ such that $A^{k}\,\tilde{\otimes}\,A^{n-k}\neq 0$. If $n=2m+1$, we obtain that $A^{k}=0$ for $k=m+1,\dots,2m+1$ and that $A^{2m+1-k}=0$ for $k=0,\dots,m$ leading to $\big(A\,\tilde{\otimes}\,I+I\,\tilde{\otimes}\,A\big)^{2m+1}=0$.
Exploiting the nilpotency of $\mathcal{K}$ now yields
@@
$\operatorname{e}^{\mathcal{K}(t-s)}\operatorname{vec}(M)=\operatorname{vec}\Big(\mathbb{E}\Big[\Phi(t,s)M\Phi(t,s)^{\top}\Big]\Big)=\mathbb{E}\Big[\Phi(t,s)\,\tilde{\otimes}\,\Phi(t,s)\Big]\operatorname{vec}(M).$
Since this identity is true for all matrices $M$, we find that $\mathbb{E}\Big[\Phi(t,s)\,\tilde{\otimes}\,\Phi(t,s)\Big]=\operatorname{e}^{\mathcal{K}(t-s)}$ and consequently
@@ 3.3 Reduced order signature approximation
We conduct a state space transformation via a nonsingular matrix $\mathcal{T}\in\mathbb{R}^{n\times n}$. To do so, we define a new state variable $\hat{\mathcal{X}}_{t}=\mathcal{T}\mathcal{X}_{t}$ and associated coefficients $(\hat{A},\hat{N}_{i},\hat{L})=(\mathcal{T}A\mathcal{T}^{-1},\mathcal{T}N_{i}\mathcal{T}^{-1},L\mathcal{T}^{-1})$. The purpose of this transformation is that an equivalent system (same quantity of interest $\mathcal{Y}$) is supposed to be obtained, in which the redundant information can be removed easily by truncation of state components. In particular, $\mathcal{T}$ is chosen in a way that it simultaneously diagonalizes $P$ and $Q$. Hence, the corresponding eigenvectors $(p_{k})$ and $(q_{k})$ are the canonical basis $(\operatorname{e}_{k})$ of $\mathbb{R}^{n}$ at the same time. Exploiting Propositions 3.4 and 3.5, we can then identify unimportant components $\langle\hat{\mathcal{X}}_{\cdot},\operatorname{e}_{k}\rangle_{\mathbb{R}^{n}}$ of $\hat{\mathcal{X}}$ and truncate those. Let us refer to the discussion below these propositions once more and provide further details below.
@@
meaning that $\mathcal{Y}$ is invariant under this transformation. Now, we see directly from (3.2) that the fundamental solution of (3.13) is $\hat{\Phi}(t,s)=\mathcal{T}\Phi(t,s)\mathcal{T}^{-1}$. Consequently, the time-averaged quadratic forms of (3.13) are
(3.14) $\hat{P}$ $=\int_{0}^{T}\mathbb{E}\left[\hat{\Phi}(u,0)\mathcal{T}z(\mathcal{T}z)^{\top}\hat{\Phi}(u,0)^{\top}\right]\mathrm{d}u=\mathcal{T}P\mathcal{T}^{\top},$
$\hat{Q}$ $=\int_{0}^{T}\mathbb{E}\left[\hat{\Phi}(u,0)^{\top}\hat{L}^{\top}\hat{L}\hat{\Phi}(u,0)\right]\mathrm{d}u=\mathcal{T}^{-\top}Q\mathcal{T}^{-1}.$
The following proposition states the particular transformation required for a simultaneous diagonalization of the time-averaged quadratic forms.
@@
Given that we have used the balancing transformation (3.15) in (3.13), we can identify less relevant directions with components of $\hat{\mathcal{X}}$ that are associated with small eigenvalues $\sigma_{k}^{2}$ of $PQ$. This is a consequence of Propositions 3.4 and 3.5, where the transformation ensures that $p_{k}=q_{k}$ is the $k$-th column of the $n\times n$ identity matrix and $\lambda_{k}=\mu_{k}=\sigma_{k}$. For that reason, the spectrum of $PQ$ delivers a good truncation criterion and therefore the intuition on how to fix the reduced dimension $\widetilde{n}\ll n$. In order to find the reduced equation, we partition the solution of (3.13) with $\mathcal{T}$ like in (3.15) as follows $\hat{\mathcal{X}}_{t}=\begin{bmatrix}\hat{\mathcal{X}}_{t}^{1}\\ \hat{\mathcal{X}}_{t}^{2}\end{bmatrix}$. $\hat{\mathcal{X}}_{t}^{1}$ taking values in $\mathbb{R}^{\widetilde{n}}$ corresponds to the large values $\sigma_{1},\ldots,\sigma_{\widetilde{n}}$ and $\hat{\mathcal{X}}_{t}^{2}$ to the small values $\sigma_{\widetilde{n}+1},\ldots,\sigma_{n}$ in the sense that $\sigma_{\widetilde{n}+1}\ll\sigma_{\widetilde{n}}$. The respective partition of the balanced coefficients